NetNews: A One-Page Guide to the Usenet

What is the Usenet/NetNews?

Usenet (Users’ Network) is a bulletin board shared among many computer systems around the world. Usenet is a logical network, sitting on top of several physical networks, including UUCP, BITNET, and the Internet. Sites on Usenet include many universities, private companies, and research organizations. Current estimates indicate that there are over two million participants at over 45,000 Usenet sites in over fifty countries, with more users and sites joining every day. Most are running the UNIX^{TM} operating system, although the number of non-UNIX sites, e.g. VMS sites and PCs, is increasing.

This Guide…

This is a one-page guide intended to provide the absolute minimum of information required to keep things working smoothly for you on the net. The best place to go for more details (the place this guide was condensed

from) is the newsgroup news.announce.newusers. Type the command rn, type g news.announce.newusers

at the prompt, and you’re ready to go. Typing h while inside rn (ReadNews) (also known as rrn, for Remote ReadNews) will give you help. The GNUS package for reading news inside of GNU Emacs and the xrn program for reading news using a windowed interface are also available. Call us for more information, or check out the Athena on-line help (see below).

Using rn or xrn

Type the command add sipb to attach the sipb locker and add it to your search path. If that doesn’t work, then there is something non-standard about your account configuration; try using these commands instead:

     attach sipb
     set path = (/mit/sipb/bin $path)

After using one of these, typing rn will bring up the text interface to the Usenet, or typing xrn will bring up the X interface.

Other Things to Check Out

You may not want to read every article in news.announce.newusers, and it’s true that not all of them are necessary. Please do read the articles entitled “Answers to Frequently Asked Questions” and “A Primer on How to Work With the Usenet Community.” You can get a lot of information about the news reader you decide to use by typing man rn or man xrn.

You might also want to check out news.newusers.questions; in particular, the posting entitled “Welcome to news.newusers.questions!” is extremely valuable for new Usenet users.

Finally, the Athena on-line help contains more detailed information about using NetNews. To access it, run help and see the “Network News” menu underneath the “Communicating with Other Users” top-level menu.

Getting Questions Answered

First of all, please do NOT ask NetNews-related questions on OLC. It is possible that you will be connected to a consultant who knows something about NetNews, but it is more likely that a busy consultant will simply direct you to us. Athena consultants are not required to know anything about SIPB software, NetNews included. If you have a question or bug report about a particular Usenet program in the sipb locker, send mail to bug-sipb@athena.mit.edu. If you have a more general question about the Usenet, send mail to usenet@athena.mit.edu. You can also call the SIPB office (3-7788) or visit us in W20-557 — we’re open nearly around the clock, every day, and we’re always happy to answer questions!

NetNews is supported by the SIPB, not by Athena. Please don’t hesitate to point out to us any problems or concerns you may have. This is for everyone to enjoy!